OpenStudy (anonymous):

Seven times a number minus the number is -48. Find the number.

OpenStudy (anonymous):

@ParthKohli can u pls help me with this?

OpenStudy (chasepit):

ok so set it up. 7*x-x=-48 you"re trying to find x so im thinking its in between 10 and 15

OpenStudy (chasepit):

actually i think it may be 8. let me plug it in

OpenStudy (anonymous):

Alright @Chasepit . Appreciate it. :)

OpenStudy (chasepit):

The answer would be negative though 7*-8-(-)8=-48 -56+8=-48 -48=-48

OpenStudy (chasepit):

you're welcome

OpenStudy (anonymous):

how you got 8?

OpenStudy (chasepit):

i just started plugging numbers in

OpenStudy (anonymous):

ohh ok. got chu. Thanks anyway. :)

OpenStudy (chasepit):

welcome